# Modified Langevin noise formalism for multiple quantum emitters in dispersive electromagnetic environments

### v1

@@ III Bright modes of the electromagnetic environment and reduced Hamiltonian
We now represent the vector field operators $\{\hat{\mathbf{f}}_{\omega}(\mathbf{r})\}$ and the scalar field operators $\{\hat{g}_{\omega\mathbf{n}\nu}\}$ through the emitter-centered modes (e.g. [18], [19], [1]). Following Ref. [19], we aim at eliminating the degrees of freedom of the medium-assisted reservoir and the scattering-assisted reservoir that are not excited by the quantum emitters. Hence, starting from Eqs. 2, 3, and 4, we introduce the auxiliary interaction operator $\hat{F}_{j}^{(M)}$ describing the coupling between the $j$-th emitter and the the medium-assisted reservoir
$\hat{F}_{j}^{(M)}=\int_{0}^{\infty}d\omega\int_{V}d^{3}\mathbf{r}\,[\mathbb{u}_{j}\cdot\mathcal{G}_{e}(\mathbf{r}_{j},\mathbf{r},\omega)\cdot\hat{\mathbf{f}}_{\omega}(\mathbf{r}_{j})+h.c.].$ (17)
@@
$\hat{F}_{j}^{(S)}=\int_{0}^{\infty}d\omega\oint d{o}_{\mathbb{n}}\,[\sum_{\nu}\,\mathbb{u}_{j}\cdot\mathbf{E}_{\omega\mathbf{n}\nu}(\mathbf{r}_{j})\hat{g}_{\omega\mathbf{n}\nu}+h.c.]\,.$ (18)
@@ III.1 Medium-Assisted Reservoir
$\hat{A}_{j}(\omega)=\int_{V}d^{3}\mathbf{r}\,\bm{\alpha}_{j}(\mathbf{r};{\omega})\cdot\hat{\mathbf{f}}_{\omega}{(\mathbf{r})}$ (21)
with $j=1,N$, where
$\bm{\alpha}_{j}(\mathbf{r};{\omega})=\mathbb{u}_{j}\cdot\mathcal{G}_{e}(\mathbf{r}_{j},\mathbf{r};\omega){,}$ (22)
@@
${\left[\hat{A}_{i}(\omega),\hat{A}_{j}^{\dagger}(\omega^{\prime})\right]}=M_{ij}(\omega)\delta\left(\omega-\omega^{\prime}\right)$ (24)
@@
$M_{ij}(\omega)=\int_{V}d^{3}\mathbf{r}\,\bm{\alpha}_{i}(\mathbf{r};{\omega})\cdot\bm{\alpha}_{j}^{*}(\mathbf{r};{\omega}).$ (25)
The $N\times N$ matrix $M=[M_{ij}]$ is a complex Gram matrix associated with the inner product $\int_{V}d\mathbf{r}^{3}\mathbb{a}(\mathbf{r})\cdot\mathbb{b}^{*}(\mathbf{r})$. By construction $M$ is Hermitian. The vector fields $\{\bm{\alpha}_{j}(\mathbf{r};\omega)\}$ are linearly independent, and then $M$ is positive definite. Since these fields are not mutually orthogonal with respect to the inner product above, $M$ is, in general, dense. Using 22, we obtain
$M_{ij}(\omega)=\int_{V}{d^{3}\mathbf{r}\,\mathbb{u}_{i}\cdot\mathcal{G}_{e}(\mathbf{r}_{i},\mathbf{r}};\omega)\mathcal{G}_{e}^{*T}(\mathbf{r}_{j},\mathbf{r};\omega)\cdot\mathbb{u}_{j}.$ (26)
We now seek a representation of $\hat{H}_{E}^{(M)}$ in terms of the scalar operators $\{\hat{A}_{j}(\omega)\}$. The vector fields $\{\bm{\alpha}_{j}(\mathbf{r};\omega)\}$ are, in general, not orthogonal with respect to the above inner product; consequently, the basis $\{\hat{A}_{j}(\omega)\}$ does not diagonalize $\hat{H}_{E}^{(M)}$. To obtain a diagonal form, we generate the orthonormal set of vector fields $\{\bm{\chi}_{i}(\mathbf{r};\omega)\}$ by linear transformation
$\bm{\chi}_{i}(\mathbf{r};\omega)=\sum_{j=1}^{N}U_{ij}(\omega)\bm{\alpha}_{j}(\mathbf{r};\omega)$ (27)
where the $N\times N$ matrix $U=[U_{ij}]$ is such that $UMU^{\dagger}=I$ and II is the identity matrix. Completing $\{\bm{\chi}_{i}(\mathbf{r};\omega)\}$ with an orthonormal set $\{\bm{\chi}^{\mathrm{dark}}_{m}(\mathbf{r})\}$, we express the field operator $\hat{\mathbf{f}}_{\omega}(\mathbf{r})$ as
$\hat{\mathbf{f}}_{\omega}(\mathbf{r})=\sum_{j=1}^{N}\bm{\chi}_{j}^{*}(\mathbf{r};\omega)\hat{C}_{j}(\omega)+\sum_{m}[\bm{\chi}_{m}^{dark}(\mathbf{r})]^{*}\hat{C}_{m}^{dark}(\omega),$ (28)
@@
Note that every vector field $\bm{\chi}_{m}^{dark}(\mathbf{r})$ does not couple to the emitters: the vector fields $\{\bm{\chi}_{i}(\mathbf{r};\omega)\}$ are the bright modes of the medium assisted field, and the vector fields $\{\bm{\chi}_{m}^{dark}(\mathbf{r})\}$ are the dark modes. The scalar bosonic operators $\hat{C}_{i}(\omega)$ are the medium-assisted bright bosonic operators of the system and $\hat{C}_{m}^{dark}(\omega)$ are the dark bosonic operators. The matrix $U$ can be chosen in various ways [19]. For our purposes, it is convenient to choose $U=M^{-1/2}$.
@@
$\hat{H}_{E}^{(M)}=\sum_{j=1}^{N}\int_{0}^{\infty}d\omega\hbar\omega\hat{C}_{j}^{\dagger}(\omega)\hat{C}_{j}(\omega)\\ +\sum_{m}\int_{0}^{\infty}d\omega\hbar\omega\sum_{m}\hat{C}_{m}^{dark\dagger}({\omega})\hat{C}_{m}^{dark}({\omega}).$ (33)
@@ III.2 Scattering-Assisted Reservoir
$\hat{B}_{j}(\omega)=\oint do_{\mathbf{n}}\sum_{\nu}\,{\beta}_{j}(\mathbf{n},\nu;\omega)\hat{g}_{\omega\mathbf{n}\nu}$ (34)
with $j=1,N$, where
${\beta}_{j}(\mathbf{n},\nu;\omega)=\mathbb{u}_{j}\cdot\mathbf{E}_{\omega\mathbf{n}\nu}(\mathbf{r}_{j}),$ (35)
@@
${\left[\hat{B}_{i}(\omega),\hat{B}_{j}^{\dagger}(\omega^{\prime})\right]}=S_{ij}(\omega)\delta\left(\omega-\omega^{\prime}\right)$ (37)
@@
$S_{ij}(\omega)=\oint do_{\mathbf{n}}\,\mathbf{u}_{i}\cdot\sum_{\nu}\mathbf{E}_{\omega\mathbf{n}\nu}(\mathbf{r}_{i})\mathbf{E}_{\omega\mathbf{n}\nu}^{*}(\mathbf{r}_{j})\cdot\mathbb{u}_{j}.$ (38)
The $N\times N$ matrix $S=[S_{ij}]$ is also a complex Gram matrix with respect to the scalar product $\int do_{\mathbf{n}}\sum_{\nu}{f}(\mathbf{n},\nu){g}^{*}(\mathbf{n},\nu)$. Therefore, $S$ is Hermitian by construction; since the vectors $\{\beta_{j}(\mathbf{n},\nu,\omega)\}$ are linearly independent, it is also positive definite. Proceeding as in the medium-assisted case, we introduce the set of functions $\left\{\xi_{j}(\mathbb{n},\nu;\omega)\right\}$, orthonormal with respect to the above scalar product. They are given by
$\xi_{i}(\mathbb{n},\nu;\omega)=\sum_{j=1}^{N}V_{ij}{\beta}_{j}(\mathbf{n},\nu;\omega)$ (39)
where the $N\times N$ matrix $V=[V_{ij}]$ is such that $VSV^{\dagger}=I$. Regarding the bright modes of the field assisted by the medium, we choose $V=S^{-1/2}$. Completing the set $\{{\xi}_{i}(\mathbb{n},\nu;\omega)\}$ with an orthonormal set $\{{\xi}^{\mathrm{dark}}_{m}\}$, we express the field operator $\hat{g}_{\omega\mathbf{n}\nu}$ as
$\hat{g}_{\omega\mathbf{n}\nu}=\sum_{j=1}^{N}\xi_{j}^{*}(\mathbb{n},\nu,\omega)\hat{D}_{j}(\omega)+\sum_{m}[{\xi}_{m}^{dark}]^{*}\hat{D}_{m}^{dark}(\omega),$ (40)
@@
As for the vector fields $\{\bm{\chi}_{m}^{dark}\}$, the scalar fields $\{\xi_{m}^{dark}\}$ do not pair up with the emitters; $\{\xi_{i}\}$ are the bright modes of the scattering-assisted field and $\{\hat{D}_{i}\}$ are the corresponding bright operators. The scalar bosonic operators $\{\hat{B}_{i}(\omega)\}$ are related to $\{\hat{D}_{i}(\omega)\}$ through the equations
@@
$\hat{H}_{E}^{(S)}=\sum_{j=1}^{N}\int_{0}^{\infty}d\omega\hbar\omega\hat{D}_{j}^{\dagger}(\omega)\hat{D}_{j}(\omega)\\ +\sum_{m}\int_{0}^{\infty}d\omega\hbar\omega\sum_{m}\hat{D}_{m}^{dark\dagger}({\omega})\hat{D}_{m}^{dark}({\omega}).$ (44)
@@ III.3 Reduced Hamiltonian
The medium-assisted and scattering-assisted dark modes are decoupled from the rest of the system, they do not affect dynamics of the quantum emitters and can be dropped from the expressions of $\hat{H}_{E}^{(M)}$ and $\hat{H}_{E}^{(S)}$. The overall system, quantum emitters + bright modes of the electromagnetic environment, is described by the reduced Hamiltonian
$\hat{H}_{red}=\hat{H}_{A}+\hat{H}_{E}^{bright}+\hat{H}_{I}$ (45)
@@
$\hat{H}_{E}^{bright}=\sum_{i=1}^{N}\int_{0}^{\infty}d\omega\hbar\omega\left[\hat{C}_{i}^{\dagger}(\omega)\hat{C}_{i}(\omega)+\hat{D}_{i}^{\dagger}(\omega)\hat{D}_{i}(\omega)\right]$ (46)
@@
$\hat{H}_{I}=\sum_{i,j=1}^{N}\hat{\mu}_{i}\left[\int_{0}^{\infty}d\omega G_{ij}^{(M)}(\omega)\hat{C}_{j}(\omega)+h.c.\right]+\\ \sum_{i,j=1}^{N}\hat{\mu}_{i}\left[\int_{0}^{\infty}d\omega G_{ij}^{(S)}(\omega)\hat{D}_{j}(\omega)+h.c.\right].$ (47)
In summary, the set of quantum emitters behaves as an open quantum system coupled to two independent reservoirs: the medium-assisted reservoir and the scattering-assisted reservoir. Each reservoir consists of $N$ independent set of bosonic modes. The medium-assisted reservoir is characterized by the coupling matrix ${G}^{(M)}(\omega)=M(\omega)^{1/2}$, and the scattering-assisted reservoir by ${G}^{(M)}(\omega)=S(\omega)^{1/2}$. In Appendix A, we provide the expression of the electric field operator due to the bright modes.
@@ IV Spectral density matrices
The matrices $G^{(M)}(\omega)$ and $G^{(S)}(\omega)$ fully characterize the interaction of the quantum emitters with the electromagnetic environment, including their mutual interactions. It is convenient to recast these coupling matrices in terms of two matrices that generalize the concepts of medium-assisted spectral density and scattering-assisted spectral density, as originally introduced in [1] and [2], for a single quantum emitter, to configurations of multiple quantum emitters.
@@
$\mathrm{\mathcal{J}}_{ij}^{(M)}(\omega)$ $=\frac{\mu_{i}\mu_{j}}{\hbar^{2}}M_{ij},$ (51a)
$\mathrm{\mathcal{J}}_{ij}^{(S)}(\omega)$ $=\frac{\mu_{i}\mu_{j}}{\hbar^{2}}S_{ij},$ (51b)
@@
$\mathrm{\mathcal{J}}_{ij}^{(M)}(\omega)$ $=\frac{\mu_{i}\mu_{j}}{\hbar^{2}}\int_{V}{d^{3}\mathbf{r}\,\mathbb{u}_{i}\cdot\mathcal{G}_{eq}(\mathbf{r}_{i},\mathbf{r}})\mathcal{G}_{eq}^{*T}(\mathbf{r}_{j},\mathbf{r})\cdot\mathbb{u}_{j},$ (52a)
$\mathrm{\mathcal{J}}_{ij}^{(S)}(\omega)$ $=\frac{\mu_{i}\mu_{j}}{\hbar^{2}}\oint do_{\mathbf{n}}\,\sum_{\nu}\mathbf{u}_{i}\cdot\mathbf{E}_{\omega\mathbf{n}\nu}(\mathbf{r}_{i})\mathbf{E}_{\omega\mathbf{n}\nu}^{*}(\mathbf{r}_{j})\cdot\mathbb{u}_{j}.$ (52b)
Both spectral density matrices are complex, Hermitian, and positive definite. The elements $\mathrm{\mathcal{J}}_{ij}^{(M)}(\omega)$ and $\mathrm{\mathcal{J}}_{ij}^{(S)}(\omega)$ are not independent; in fact, as a consequence of 14
$\mathrm{\mathcal{J}}_{ij}^{(M)}(\omega)+\mathrm{\mathcal{J}}_{ij}^{(S)}(\omega)=\frac{1}{2\pi}\Gamma_{ij}(\omega)$ (53)
@@
$\Gamma_{ij}(\omega)=\frac{2\omega^{2}}{\hbar\varepsilon_{0}c^{2}}(\mu_{i}\mathbb{u}_{i})\cdot\operatorname{Im}\left[\mathcal{G}_{\omega}\left(\mathbf{r}_{i}.\mathbf{r}_{j}\right)\right]\cdot(\mu_{j}\mathbb{u}_{j}).$ (54)
@@
The medium-assisted spectral density and the scattering-assisted spectral density have a direct interpretation within classical electrodynamics. Consider the classical electromagnetic field generated, in the presence of dielectric objects, by $N$ electric dipoles with dipole moments $\mu_{j}\mathbf{u}_{j}$, located at ${\mathbf{r}}_{j}$ with $j=1,N$ and oscillating at frequency $\omega$. We find that $\frac{\pi}{2}\hbar\omega\sum_{i,j=1}^{N}\mathcal{J}^{(M)}_{ij}(\omega)$ is equal to the average electromagnetic power absorbed by the dielectric bodies; $\frac{1}{4}(\hbar\omega)\sum_{i,j=1}^{N}\Gamma_{ij}(\omega)$ is equal to the averaged electromagnetic power emitted by the set of electric dipoles. Then, $\frac{\pi}{2}\hbar\omega\sum_{i,j=1}^{N}\mathcal{J}^{(S)}_{ij}(\omega)$ is equal to the average electromagnetic power radiated toward infinity. Accordingly, Eq. 53 represents the statement of the Poynting theorem for the investigated scenario. See Appendix C for a full discussion.
V Correlator matrix and surrogate environment
The dynamic of $N$ quantum emitters coupled to $N$ medium-assisted bosonic modes characterized by the spectral density matrix $\mathcal{J}^{(M)}(\omega)$ and to $N$ scattering-assisted bosonic modes characterized by the spectral density matrix $\mathcal{J}^{(S)}(\omega)$ can, in principle, be computed numerically with a variety of methods [20, 21]. However, this is often computationally demanding. Under specific conditions, analogous to the single quantum emitter [1], [2] case, the actions of the $2N$ bosonic modes can be represented by $N$ bosonic modes characterized by an equivalent $N\times N$ spectral density matrix, in the same spirit of [22, 23].
We assume that: (i) the initial quantum state of the entire system is a product state, that is, $\hat{\rho}(0)=\hat{\rho}_{A}(0)\otimes\hat{\rho}_{E}(0)$, where $\hat{\rho}(0)$, $\hat{\rho}_{A}(0)$ and $\hat{\rho}_{E}(0)$ are the density operators of the entire system, of the set of quantum emitters and of the electromagnetic environment, respectively; (ii) the initial state of the electromagnetic environment is Gaussian. Then, the evolution of the reduced density operator of the set of quantum emitters $\hat{\rho}_{A}(t)=\text{Tr}_{E}\left[\rho(t)\right]$ depends only on the expectation values ${F}_{j}(t)$ and the two-time correlation functions ${C}_{ij}(t+\tau;t)$ (e.g. [24]) of the environment interaction operators $\hat{F}_{j}$ with i,j=1,2,,Ni,j=1,2,...,N,

### v2

@@ III Bright modes of the electromagnetic environment and reduced Hamiltonian
We now represent the vector field operators $\{\hat{\mathbf{f}}_{\omega}(\mathbf{r})\}$ and the scalar field operators $\{\hat{g}_{\omega\mathbf{n}\nu}\}$ in terms of the emitter-centered modes (e.g. [24], [25], [1]). Following Ref. [25], we aim to eliminate the degrees of freedom of the medium-assisted reservoir and the scattering-assisted reservoir that are not excited by the quantum emitters. Hence, starting from Eqs. 2, 3, and 4, we introduce the auxiliary interaction operator $\hat{F}_{j}^{(M)}$ describing the coupling between the $j$-th emitter and the medium-assisted reservoir
$\hat{F}_{j}^{(M)}=\int_{0}^{\infty}d\omega\int_{V}d^{3}\mathbf{r}\,[\mathbf{u}_{j}\cdot\mathcal{G}_{e}(\mathbf{r}_{j},\mathbf{r},\omega)\cdot\hat{\mathbf{f}}_{\omega}(\mathbf{r})+h.c.].$ (17)
@@
$\hat{F}_{j}^{(S)}=\int_{0}^{\infty}d\omega\oint d{o}_{\mathbf{n}}\,[\sum_{\nu}\,\mathbf{u}_{j}\cdot\mathbf{E}_{\omega\mathbf{n}\nu}(\mathbf{r}_{j})\hat{g}_{\omega\mathbf{n}\nu}+h.c.]\,.$ (18)
@@ III.1 Medium-Assisted Reservoir
$\hat{A}_{j}(\omega)=\int_{V}d^{3}\mathbf{r}\,\boldsymbol{\alpha}_{j}(\mathbf{r};{\omega})\cdot\hat{\mathbf{f}}_{\omega}{(\mathbf{r})},$ (21)
with $j=1,\ldots,N$, where
$\boldsymbol{\alpha}_{j}(\mathbf{r};{\omega})=\mathbf{u}_{j}\cdot\mathcal{G}_{e}(\mathbf{r}_{j},\mathbf{r};\omega){,}$ (22)
@@
${\left[\hat{A}_{i}(\omega),\hat{A}_{j}^{\dagger}(\omega^{\prime})\right]}=M_{ij}(\omega)\delta\left(\omega-\omega^{\prime}\right),$ (24)
@@
$M_{ij}(\omega)=\int_{V}d^{3}\mathbf{r}\,\boldsymbol{\alpha}_{i}(\mathbf{r};{\omega})\cdot\boldsymbol{\alpha}_{j}^{*}(\mathbf{r};{\omega}).$ (25)
The $N\times N$ matrix $M=[M_{ij}]$ is a complex Gram matrix associated with the inner product $\int_{V}d\mathbf{r}^{3}\mathbf{a}(\mathbf{r})\cdot\mathbf{b}^{*}(\mathbf{r})$. By construction $M$ is Hermitian. The vector fields $\{\boldsymbol{\alpha}_{j}(\mathbf{r};\omega)\}$ are linearly independent, and then $M$ is positive definite. Since these fields are not mutually orthogonal with respect to the inner product above, $M$ is, in general, dense. Using 22, we obtain
$M_{ij}(\omega)=\int_{V}d^{3}\mathbf{r}\,\mathbf{u}_{i}\cdot\mathcal{G}_{e}(\mathbf{r}_{i},\mathbf{r};\omega)\mathcal{G}_{e}^{*T}(\mathbf{r}_{j},\mathbf{r};\omega)\cdot\mathbf{u}_{j}.$ (26)
We now seek a representation of $\hat{H}_{E}^{(M)}$ in terms of the scalar operators $\{\hat{A}_{j}(\omega)\}$. The vector fields $\{\boldsymbol{\alpha}_{j}(\mathbf{r};\omega)\}$ are, in general, not orthogonal with respect to the above inner product; consequently, the basis $\{\hat{A}_{j}(\omega)\}$ does not diagonalize $\hat{H}_{E}^{(M)}$. To obtain a diagonal form, we generate the orthonormal set of vector fields $\{\boldsymbol{\chi}_{i}(\mathbf{r};\omega)\}$ by applying the linear transformation
$\boldsymbol{\chi}_{i}(\mathbf{r};\omega)=\sum_{j=1}^{N}U_{ij}(\omega)\boldsymbol{\alpha}_{j}(\mathbf{r};\omega),$ (27)
where the $N\times N$ matrix $U=[U_{ij}]$ is such that $UMU^{\dagger}=I$ and II is the identity matrix. Completing $\{\boldsymbol{\chi}_{i}(\mathbf{r};\omega)\}$ with an orthonormal set $\{\boldsymbol{\chi}^{\mathrm{dark}}_{m}(\mathbf{r};\omega)\}$, we express the field operator $\hat{\mathbf{f}}_{\omega}(\mathbf{r})$ as
$\hat{\mathbf{f}}_{\omega}(\mathbf{r})=\sum_{j=1}^{N}\boldsymbol{\chi}_{j}^{*}(\mathbf{r};\omega)\hat{C}_{j}(\omega)+\sum_{m}[\boldsymbol{\chi}_{m}^{\text{dark}}(\mathbf{r};\omega)]^{*}\hat{C}_{m}^{\text{dark}}(\omega),$ (28)
@@
Note that every vector field $\boldsymbol{\chi}_{m}^{\text{dark}}(\mathbf{r};\omega)$ does not couple to the emitters: the vector fields $\{\boldsymbol{\chi}_{i}(\mathbf{r};\omega)\}$ are the bright modes of the medium assisted field, while the vector fields $\{\boldsymbol{\chi}_{m}^{\text{dark}}(\mathbf{r};\omega)\}$ are the dark modes. The scalar bosonic operators $\hat{C}_{i}(\omega)$ are the medium-assisted bright bosonic operators of the system and $\hat{C}_{m}^{\text{dark}}(\omega)$ are the dark bosonic operators. The matrix $U$ can be chosen in various ways [25]. For our purposes, it is convenient to choose $U=M^{-1/2}$.
@@
$\hat{H}_{E}^{(M)}=\sum_{j=1}^{N}\int_{0}^{\infty}d\omega\hbar\omega\hat{C}_{j}^{\dagger}(\omega)\hat{C}_{j}(\omega)\\ +\sum_{m}\int_{0}^{\infty}d\omega\hbar\omega\hat{C}_{m}^{dark\dagger}({\omega})\hat{C}_{m}^{\text{dark}}({\omega}).$ (33)
@@ III.2 Scattering-Assisted Reservoir
$\hat{B}_{j}(\omega)=\oint do_{\mathbf{n}}\sum_{\nu}\,{\beta}_{j}(\mathbf{n},\nu;\omega)\hat{g}_{\omega\mathbf{n}\nu},$ (34)
with $j=1,\ldots,N$, where
${\beta}_{j}(\mathbf{n},\nu;\omega)=\mathbf{u}_{j}\cdot\mathbf{E}_{\omega\mathbf{n}\nu}(\mathbf{r}_{j}),$ (35)
@@
${\left[\hat{B}_{i}(\omega),\hat{B}_{j}^{\dagger}(\omega^{\prime})\right]}=S_{ij}(\omega)\delta\left(\omega-\omega^{\prime}\right),$ (37)
@@
$S_{ij}(\omega)=\oint do_{\mathbf{n}}\,\mathbf{u}_{i}\cdot\sum_{\nu}\mathbf{E}_{\omega\mathbf{n}\nu}(\mathbf{r}_{i})\mathbf{E}_{\omega\mathbf{n}\nu}^{*}(\mathbf{r}_{j})\cdot\mathbf{u}_{j}.$ (38)
The $N\times N$ matrix $S=[S_{ij}]$ is also a complex Gram matrix with respect to the scalar product $\int do_{\mathbf{n}}\sum_{\nu}{f}(\mathbf{n},\nu){g}^{*}(\mathbf{n},\nu)$. Therefore, $S$ is Hermitian by construction; since the vectors $\{\beta_{j}(\mathbf{n},\nu,\omega)\}$ are linearly independent, it is also positive definite. Proceeding as in the medium-assisted case, we introduce the set of functions $\left\{\xi_{j}(\mathbf{n},\nu;\omega)\right\}$, orthonormal with respect to the above scalar product. They are given by
$\xi_{i}(\mathbf{n},\nu;\omega)=\sum_{j=1}^{N}V_{ij}{\beta}_{j}(\mathbf{n},\nu;\omega),$ (39)
where the $N\times N$ matrix $V=[V_{ij}]$ is such that $VSV^{\dagger}=I$. In analogy with the medium-assisted case, we choose $V=S^{-1/2}$. Completing the set $\{{\xi}_{i}(\mathbf{n},\nu;\omega)\}$ with an orthonormal set $\{{\xi}^{\mathrm{dark}}_{m}\}$, we express the field operator $\hat{g}_{\omega\mathbf{n}\nu}$ as
$\hat{g}_{\omega\mathbf{n}\nu}=\sum_{j=1}^{N}\xi_{j}^{*}(\mathbf{n},\nu,\omega)\hat{D}_{j}(\omega)+\sum_{m}[{\xi}_{m}^{\text{dark}}]^{*}\hat{D}_{m}^{\text{dark}}(\omega),$ (40)
@@
As for the medium assisted case, the scalar fields $\{\xi_{m}^{\text{dark}}\}$ do not couple to the emitters; $\{\xi_{i}\}$ are the bright modes of the scattering-assisted field and $\{\hat{D}_{i}\}$ are the corresponding bright operators. The scalar bosonic operators $\{\hat{B}_{i}(\omega)\}$ are related to $\{\hat{D}_{i}(\omega)\}$ through the equations
@@
$\hat{H}_{E}^{(S)}=\sum_{j=1}^{N}\int_{0}^{\infty}d\omega\hbar\omega\hat{D}_{j}^{\dagger}(\omega)\hat{D}_{j}(\omega)\\ +\sum_{m}\int_{0}^{\infty}d\omega\hbar\omega\hat{D}_{m}^{dark\dagger}({\omega})\hat{D}_{m}^{\text{dark}}({\omega}).$ (44)
@@ III.3 Reduced Hamiltonian
The medium-assisted and scattering-assisted dark modes are decoupled from the rest of the system, do not affect the dynamics of the quantum emitters, and can be dropped from the expressions of $\hat{H}_{E}^{(M)}$ and $\hat{H}_{E}^{(S)}$. The overall system, i.e., quantum emitters + bright modes of the electromagnetic environment, is described by the reduced Hamiltonian
$\hat{H}_{red}=\hat{H}_{A}+\hat{H}_{E}^{\text{bright}}+\hat{H}_{I},$ (45)
@@
$\hat{H}_{E}^{\text{bright}}=\sum_{i=1}^{N}\int_{0}^{\infty}d\omega\hbar\omega\left[\hat{C}_{i}^{\dagger}(\omega)\hat{C}_{i}(\omega)+\hat{D}_{i}^{\dagger}(\omega)\hat{D}_{i}(\omega)\right],$ (46)
@@
$\hat{H}_{I}=-\sum_{i,j=1}^{N}\hat{\mu}_{i}\left[\int_{0}^{\infty}d\omega G_{ij}^{(M)}(\omega)\hat{C}_{j}(\omega)+h.c.\right]\\ -\sum_{i,j=1}^{N}\hat{\mu}_{i}\left[\int_{0}^{\infty}d\omega G_{ij}^{(S)}(\omega)\hat{D}_{j}(\omega)+h.c.\right].$ (47)
In summary, the set of quantum emitters behaves as an open quantum system coupled to two independent reservoirs: the medium-assisted reservoir and the scattering-assisted reservoir. Each reservoir consists of $N$ independent set of bosonic modes. The medium-assisted reservoir is characterized by the coupling matrix ${G}^{(M)}(\omega)=M(\omega)^{1/2}$, and the scattering-assisted reservoir by ${G}^{(S)}(\omega)=S(\omega)^{1/2}$. In Appendix A, we provide the expression of the electric field operator due to the bright modes.
@@ IV Spectral density matrices
The matrices $G^{(M)}(\omega)$ and $G^{(S)}(\omega)$ fully characterize the interaction of the quantum emitters with the electromagnetic environment, including their mutual interactions. It is convenient to recast these coupling matrices in terms of two matrices that generalize the concepts of medium-assisted spectral density and scattering-assisted spectral density, as originally introduced in [1] and [2] for a single quantum emitter, to configurations of multiple quantum emitters.
@@
$\mathcal{J}_{ij}^{(M)}(\omega)$ $=\frac{\mu_{i}\mu_{j}}{\hbar^{2}}M_{ij},$ (51a)
$\mathcal{J}_{ij}^{(S)}(\omega)$ $=\frac{\mu_{i}\mu_{j}}{\hbar^{2}}S_{ij},$ (51b)
@@
$\mathcal{J}_{ij}^{(M)}(\omega)$ $=\frac{\mu_{i}\mu_{j}}{\hbar^{2}}\int_{V}d^{3}\mathbf{r}\,\mathbf{u}_{i}\cdot\mathcal{G}_{e}(\mathbf{r}_{i},\mathbf{r})\mathcal{G}_{e}^{*T}(\mathbf{r}_{j},\mathbf{r})\cdot\mathbf{u}_{j},$ (52a)
$\mathcal{J}_{ij}^{(S)}(\omega)$ $=\frac{\mu_{i}\mu_{j}}{\hbar^{2}}\oint do_{\mathbf{n}}\,\sum_{\nu}\mathbf{u}_{i}\cdot\mathbf{E}_{\omega\mathbf{n}\nu}(\mathbf{r}_{i})\mathbf{E}_{\omega\mathbf{n}\nu}^{*}(\mathbf{r}_{j})\cdot\mathbf{u}_{j}.$ (52b)
Both spectral density matrices are complex, Hermitian, and positive definite. The elements $\mathcal{J}_{ij}^{(M)}(\omega)$ and $\mathcal{J}_{ij}^{(S)}(\omega)$ are not independent; in fact, as a consequence of 14
$\mathcal{J}_{ij}^{(M)}(\omega)+\mathcal{J}_{ij}^{(S)}(\omega)=\frac{1}{2\pi}\Gamma_{ij}(\omega),$ (53)
@@
$\Gamma_{ij}(\omega)=\frac{2\omega^{2}}{\hbar\varepsilon_{0}c^{2}}(\mu_{i}\mathbf{u}_{i})\cdot\operatorname{Im}\left[\mathcal{G}_{\omega}\left(\mathbf{r}_{i},\mathbf{r}_{j}\right)\right]\cdot(\mu_{j}\mathbf{u}_{j}).$ (54)
@@
The medium-assisted spectral density and the scattering-assisted spectral density have a direct interpretation within classical electrodynamics. Consider the classical electromagnetic field generated, in the presence of dielectric objects, by $N$ electric dipoles with dipole moments $\mu_{j}\mathbf{u}_{j}$, located at ${\mathbf{r}}_{j}$ with $j=1,\ldots,N$ and oscillating at frequency $\omega$. We find that $\frac{\pi}{2}\hbar\omega\sum_{i,j=1}^{N}\mathcal{J}^{(M)}_{ij}(\omega)$ is equal to the average electromagnetic power absorbed by the dielectric bodies; $\frac{1}{4}(\hbar\omega)\sum_{i,j=1}^{N}\Gamma_{ij}(\omega)$ is equal to the averaged electromagnetic power emitted by the set of electric dipoles. Then $\frac{\pi}{2}\hbar\omega\sum_{i,j=1}^{N}\mathcal{J}^{(S)}_{ij}(\omega)$ is equal to the average electromagnetic power radiated toward infinity. Accordingly, Eq. 53 represents the statement of the Poynting theorem for the investigated scenario. See Appendix C for a full discussion.
V Correlator matrices and surrogate environment
The dynamics of $N$ quantum emitters coupled to $N$ medium-assisted bosonic modes characterized by the spectral density matrix $\mathcal{J}^{(M)}(\omega)$ and to $N$ scattering-assisted bosonic modes characterized by the spectral density matrix $\mathcal{J}^{(S)}(\omega)$ can, in principle, be computed numerically with a variety of methods [26, 27]. However, this is often computationally demanding. Under specific conditions, analogous to the single quantum emitter case [1], [2], the actions of the $2N$ bosonic modes can be represented by only $N$ bosonic modes characterized by an equivalent $N\times N$ spectral density matrix, in the same spirit as in [28, 29].
We assume that: (i) the quantum state of the composite system at the initial time $t_{0}=0$ is a product state, that is, $\hat{\rho}(0)=\hat{\rho}_{A}(0)\otimes\hat{\rho}_{E}^{(M)}(0)\otimes\hat{\rho}_{E}^{(S)}(0)$, where $\hat{\rho}(0)$, $\hat{\rho}_{A}(0)$, $\hat{\rho}_{E}^{(M)}(0)$ and $\hat{\rho}_{E}^{(S)}(0)$ are the initial density operators of the entire system, the set of quantum emitters, the medium-assisted reservoir and the scattering-assisted reservoir, respectively; (ii) the initial states of both reservoirs are Gaussian. Then, the evolution of the reduced density operator of the set of quantum emitters $\hat{\rho}_{A}(t)=\text{Tr}_{E}\left[\hat{\rho}(t)\right]$ depends only on the expectation values ${F}_{j}(t)$ and the two-time correlation functions ${C}_{ij}(t+\tau;t)$ (e.g. [30]) of the environment interaction operators $\hat{F}_{j}$ with $i,j=1,2,\ldots,N$,